# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

### v2

@@ Appendix A Mathematical Proofs
A.2 Proof of Asymptotic Coverage (Theorem 4.2)
Theorem Statement. Under the online update $C_{t+1}=C_{t}+\gamma_{t}e_{t}$, where $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$, with step-sizes $\gamma_{t}$ satisfying the Robbins-Monro conditions ($\sum\gamma_{t}=\infty,\sum\gamma_{t}^{2}<\infty$), and assuming ergodicity, then almost surely:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha.$
Proof. This proof relies on the standard convergence results for stochastic approximation algorithms, as established by Robbins and Monro (1951) and further developed by Kushner and Yin (2003).
Let $\mathcal{F}_{t-1}$ be the filtration representing all information available up to time $t-1$. The update rule for our conformal threshold, $C_{t}$, is:
$C_{t+1}=C_{t}+\gamma_{t}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha).$
This is a stochastic approximation algorithm designed to find the root of the function $f(C)=\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C]$. The algorithm seeks the value $C^{*}$ such that $f(C^{*})=0$.
(i) Martingale Decomposition.
The update can be decomposed into a deterministic component and a stochastic noise component. Let the coverage error be $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$. We can write:
$C_{t+1}-C_{t}=\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]+\gamma_{t}(e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]).$
The term $\Delta_{t}=e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is a martingale-difference sequence, as $\mathbb{E}[\Delta_{t}|\mathcal{F}_{t-1}]=0$. This term represents the zero-mean noise driving the process. The term $\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is the “signal” that pushes $C_{t}$ towards the desired equilibrium.
(ii) Almost Sure Convergence.
The Robbins-Monro theorem guarantees that $C_{t}$ converges to the unique root $C^{*}$ almost surely, provided certain conditions are met. Our setup satisfies these conditions:
  1. 1.
    The learning rate $\gamma_{t}$ satisfies $\sum_{t=1}^{\infty}\gamma_{t}=\infty$ and $\sum_{t=1}^{\infty}\gamma_{t}^{2}<\infty$, by construction.
  2. 2.
    The error term $e_{t}$ is bounded, since $e_{t}\in[-\alpha,1-\alpha]$.
  3. 3.
    Under standard assumptions of ergodicity and continuity of the underlying return distribution, there exists a unique threshold $C^{*}$ such that the expected coverage error is zero: $\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C^{*}]=0$.
Given these conditions, standard stochastic approximation theory guarantees that $C_{t}\to C^{*}$ almost surely.
(iii) Convergence of the Coverage Fraction.
As $C_{t}$ converges to the stable point $C^{*}$, the system reaches an equilibrium where the expected coverage error is zero. By Kronecker’s lemma (Hall and Hyde, 1980), a consequence of the convergence of the stochastic approximation is that the long-run average of the error term also converges to zero:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}e_{t}=0\quad a.s.$
Substituting the definition of $e_{t}$:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha)=0.$
By the linearity of limits, this directly implies that the long-run miscoverage fraction converges to $\alpha$:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=\alpha.$
The long-run fraction of observations falling inside the interval is simply one minus the miscoverage fraction. Therefore,
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=1-\alpha.$
This completes the proof.
Appendix B Appendix: Supplementary Visualizations
@@
Appendix C Appendix: Supplementary Sensitivity Analysis
This section provides the full sensitivity analysis results for BTC-USD (3 and Gold (4), complementing the main analysis on the S&P 500 presented in the results section. The findings confirm that the TCP framework is robust and performs consistently across different asset classes and hyperparameter settings.
Table 3: TCP Sensitivity Analysis on BTC-USD Data
@@ Appendix C Appendix: Supplementary Sensitivity Analysis
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8716 9.6583
100 0.010 0.8806 9.6309
100 0.050 0.9089 9.5891
252 0.005 0.8751 9.6644
252 0.010 0.8854 9.6109
252 0.050 0.9165 9.5226
500 0.005 0.8767 9.7543
500 0.010 0.8861 9.7121
500 0.050 0.9157 9.6235
Table 4: TCP Sensitivity Analysis on Gold Data
@@
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8654 2.2101
100 0.010 0.8751 2.1987
100 0.050 0.9034 2.1643
252 0.005 0.8703 2.2214
252 0.010 0.8806 2.2021
252 0.050 0.9096 2.1589
500 0.005 0.8732 2.2456
500 0.010 0.8821 2.2298
500 0.050 0.9103 2.1887

### v3

@@ Appendix A MATHEMATICAL PROOFS
A.2 Proof of Asymptotic Calibration (Theorem 4.2)
Theorem. Under (A1)–(A4) and the update $C_{t+1}=C_{t}+\gamma_{t}e_{t}$ with $e_{t}=\mathbf{1}\{r_{t}\notin[\ell_{t},u_{t}]\}-\alpha$, we have $C_{t}\to C^{\star}$ a.s. where $g(C^{\star})=1-\alpha$, and $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}\to 1-\alpha$ a.s.
Proof. Step 1 (Notation and stability). Let $\mathcal{F}_{t}$ be the filtration generated by the data up to time $t$. Define $g(C)=\Pr\!\big(r_{t}\in[\ell_{t},u_{t}]\mid C_{t}=C\big)$ and $\phi(C)=\mathbb{E}[e_{t}\mid C_{t}=C]=\alpha-g(C)$. By (A3), $\phi$ is continuous and strictly decreasing in a neighborhood of its unique root $C^{\star}$ with $\phi(C^{\star})=0$. Under (A1) the two-sided nonconformity scores are bounded, hence there exists $S<\infty$ such that valid thresholds lie in $[0,S]$. Without loss of generality we assume the recursion is kept in a compact interval (e.g., via projection onto $[0,S]$); this is standard in SA and does not change the produced intervals.
Step 2 (SA decomposition). Decompose the coverage error into signal + noise:
$e_{t}\;=\;\mathbb{E}[e_{t}\mid\mathcal{F}_{t-1}]\;+\;\Delta_{t}\;=\;\phi(C_{t})\;+\;\Delta_{t},$
where $\Delta_{t}:=e_{t}-\mathbb{E}[e_{t}\mid\mathcal{F}_{t-1}]$ is a martingale-difference sequence with $\mathbb{E}[\Delta_{t}\mid\mathcal{F}_{t-1}]=0$ and uniformly bounded second moments by (A2). The recursion becomes
$C_{t+1}\;=\;C_{t}\;+\;\gamma_{t}\{\phi(C_{t})+\Delta_{t}\}.$ (3)
Step 3 (Convergence of $C_{t}$ to $C^{\star}$). Consider the Lyapunov function $V(c)=(c-C^{\star})^{2}$. From (3),
$C_{t+1}-C^{\star}\;=\;(C_{t}-C^{\star})+\gamma_{t}\{\phi(C_{t})+\Delta_{t}\},$
hence
$V(C_{t+1})$ $=(C_{t}-C^{\star})^{2}\;+\;2\gamma_{t}(C_{t}-C^{\star})\{\phi(C_{t})+\Delta_{t}\}\;+\;\gamma_{t}^{2}\{\phi(C_{t})+\Delta_{t}\}^{2}.$
Taking conditional expectation and using $\mathbb{E}[\Delta_{t}\mid\mathcal{F}_{t-1}]=0$,
$\mathbb{E}[V(C_{t+1})\mid\mathcal{F}_{t-1}]\;=\;V(C_{t})\;+\;2\gamma_{t}(C_{t}-C^{\star})\phi(C_{t})\;+\;\gamma_{t}^{2}\,\mathbb{E}\big[\{\phi(C_{t})+\Delta_{t}\}^{2}\mid\mathcal{F}_{t-1}\big].$
By (A1)–(A2) and compactness of the state space, there exists $K<\infty$ with $\mathbb{E}[(\phi(C_{t})+\Delta_{t})^{2}\mid\mathcal{F}_{t-1}]\leq K$ a.s. By (A3), $\phi$ is strictly decreasing near $C^{\star}$, hence there exists $\eta>0$ and $\rho>0$ such that $(C-C^{\star})\phi(C)\leq-\eta(C-C^{\star})^{2}$ whenever $|C-C^{\star}|\leq\rho$. Since iterates remain in a compact interval and step-sizes decrease, standard SA stability (together with projection if used) ensures that eventually $|C_{t}-C^{\star}|\leq\rho$ a.s. (alternatively, one may argue by contradiction using the drift term). Therefore, for all large $t$,
$\mathbb{E}[V(C_{t+1})\mid\mathcal{F}_{t-1}]\;\leq\;V(C_{t})\;-\;2\eta\,\gamma_{t}(C_{t}-C^{\star})^{2}\;+\;K\,\gamma_{t}^{2}.$
By the Robbins–Siegmund supermartingale convergence lemma, since $\sum_{t}\gamma_{t}^{2}<\infty$ and $V(\cdot)\geq 0$, we obtain that $V(C_{t})$ converges a.s. and $\sum_{t}\gamma_{t}(C_{t}-C^{\star})^{2}<\infty$ a.s. Because $\sum_{t}\gamma_{t}=\infty$ (A4), the latter implies $\liminf_{t}(C_{t}-C^{\star})^{2}=0$, and together with the monotone drift near $C^{\star}$ yields $C_{t}\to C^{\star}$ almost surely.
Step 4 (Time-average calibration). Recall $e_{t}=\phi(C_{t})+\Delta_{t}$. Since $C_{t}\to C^{\star}$ and $\phi$ is continuous, $\phi(C_{t})\to 0$ a.s.; by Cesàro averaging, $\frac{1}{T}\sum_{t=1}^{T}\phi(C_{t})\to 0$ a.s. By (A2) and square-integrability of $\Delta_{t}$, the strong law for martingale differences (e.g., Hall & Heyde, 1980) implies $\frac{1}{T}\sum_{t=1}^{T}\Delta_{t}\to 0$ a.s. Hence
$\frac{1}{T}\sum_{t=1}^{T}e_{t}\;=\;\frac{1}{T}\sum_{t=1}^{T}\phi(C_{t})\;+\;\frac{1}{T}\sum_{t=1}^{T}\Delta_{t}\;\xrightarrow{\text{a.s.}}\;0.$
Equivalently, $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}\to 1-\alpha$ a.s., which completes the proof.
Appendix B SUPPLEMENTARY VISUALIZATIONS
@@
Appendix C VISUALIZATIONS OF THE TCP VS TCP-RM TRACE PLOTS
This section provides the visualizations of the TCP vs TCP-RM trace plots for BTC-USD (Figure 5) and Gold (Figure 6), complementing the main visualization for the S&P 500 presented in the results section.
Figure 5: TCP vs. TCP-RM on BTC-USD (95% intervals). Same layout as Fig. 2. Large, frequent shocks lead to sustained but controlled elevations in $C_{\text{eff}}$, keeping rolling coverage close to 95% while allowing intervals to shrink as volatility recedes.
Figure 6: TCP vs. TCP-RM on Gold (95% intervals). Same layout as Fig. 2. Spikes in $C_{\text{eff}}$ are concentrated around crisis windows; elsewhere, $C_{\text{RM}}$ stays near zero and the method behaves like standard split-conformal TCP.
Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
This section provides the full sensitivity analysis results for BTC-USD (Table 3) and Gold (Table 4), complementing the main analysis on the S&P 500 presented in the results section. The findings confirm that the TCP-RM framework is robust and performs consistently across different asset classes and hyperparameter settings.
Table 3: TCP-RM sensitivity on BTC-USD with $|\mathcal{C}|=40$.
@@ Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
100 0.005 0.9519 17.1883
100 0.010 0.9519 17.1899
100 0.050 0.9519 17.2031
252 0.005 0.9593 10.7970
252 0.010 0.9593 10.7974
252 0.050 0.9593 10.8004
500 0.005 0.9517 9.0519
500 0.010 0.9525 9.0544
500 0.050 0.9542 9.0737
Table 4: TCP-RM sensitivity on Gold with $|\mathcal{C}|=40$.
@@
Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
100 0.005 0.9538 3.7060
100 0.010 0.9532 3.7066
100 0.050 0.9532 3.7109
252 0.005 0.9517 2.6241
252 0.010 0.9524 2.6245
252 0.050 0.9524 2.6294
500 0.005 0.9534 2.4083
500 0.010 0.9609 2.4212
500 0.050 0.9542 2.4105
Appendix E SUPPLEMENTARY SENSITIVITY ANALYSIS ON TCP (NO RM)
For completeness, in Tables 5, 6 and 7, we also report sensitivity for TCP (split-conformal only, no Robbins–Monro offset). We vary $w\in\{100,252,500\}$ with $|\mathcal{C}|=40$ for each asset.
Table 5: TCP sensitivity on S&P 500 with $|\mathcal{C}|=40$.
Window Size ($w$) Coverage Interval Width
100 0.9557 4.5174
252 0.9455 3.4538
500 0.9450 3.2376
Table 6: TCP sensitivity on BTC-USD with $|\mathcal{C}|=40$.
Window Size ($w$) Coverage Interval Width
100 0.9519 17.1866
252 0.9593 10.7967
500 0.9500 9.0469
Table 7: TCP sensitivity on Gold with $|\mathcal{C}|=40$.
Window Size ($w$) Coverage Interval Width
100 0.9538 3.7055
252 0.9517 2.6234
500 0.9417 2.3872
As with TCP-RM, larger $w$ generally yields narrower intervals. Without the adaptive offset, coverage can dip slightly below $95\%$ in some settings, underscoring the benefit of the online calibration layer.